OpenStudy (ariana):

The average adult contains 5.1 red blood cells per microliters of blood. (4.7 Liters of blood is average for adult)What is the molar concentration of red blood cells in human blood? Molar concentration = moles solute / liters of solution The average diameter of a blood cell is 7.7 μm. How many meters would the blood cells from the average human span if placed end to end? (the average adult contains 5.1 blood cells)

OpenStudy (festinger):

\[5.1/\mu l \times 4.7l=5.1/10^{-6} \times 4.7 = 2.4*10^{7} \,\mathrm{red\,blood\,cells}\] \[2.4*10^{7} \times7.7\mu m=2.4*10^{7} \times 7.7*10^{-6}m = 185m\]